You are given the following information about an economy:Gross Investment $40Govt. purchases of goods & service 30GNP 200X –M -20Personal Tax 60Govt. transfer 25Interest payments from the Govt.to domestic Pvt. Sector 15Factor income received from the rest of the would 7Factor payment made to rest of would 9Calculate:a.Consumption b) GDP c) Net factor payment from abroadb.Pvt. Saving e) Public Saving.
1
Expert's answer
2016-11-14T09:59:12-0500
GI = $40, G = 30, GNP = 200, NX = X – M = 20, T = 60, Govt. transfer = 25, Interest payments from the Govt.to domestic Pvt.Sector = 15, Factor income received from the rest of the world FI = 7, Factor payment made to rest of world FP = 9.
a) Consumption C = GDP - NX - GI - G = 202 - 20 - 40 - 30 = 112.
b) GDP = GNP + FP - FI = 200 + 9 - 7 = 202.
c) Net factor payment from abroad NFP = FP - FI = 9 - 7 = 2.
d) Pvt. Saving PrS = C - T = 112 - 60 = 42.
e) Public Saving PS = GI = 40.